40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 19 23:50:44 2013, Jenne, Update, LSC, PRCL locking again 
    Reply  Thu Jun 20 01:13:18 2013, Jenne, Update, LSC, PRCL locking again - no ASC success 
       Reply  Wed Jun 26 04:38:02 2013, Jenne, Update, LSC, PRCL locking again - ASC success PRMI_ASC_yawOnly_powerSpectra_25June2013.pdfPRMI_ASC_yawOnly_25June2013_mag.pdfPRMI_ASC_yawOnly_25June2013_phase.pdfPRMI_ASC_yawOnly_25June2013_coherence.pdfSensMatMeas_26June2013.png
          Reply  Thu Jun 27 17:09:41 2013, Jenne, Update, LSC, PRCL locking again - POP PIT work MeasuredVsModeledASCyaw.png
             Reply  Fri Jun 28 02:12:41 2013, Jenne, Update, LSC, PRCL locking again - ASC work PRMI_ASC_powerSpectra_27June2013.pdfScreenshot-Untitled_Window.png
Message ID: 8731     Entry time: Thu Jun 20 01:13:18 2013     In reply to: 8730     Reply to this: 8753
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRCL locking again - no ASC success 

I didn't have any success with the ASC tonight.  I copied over the filters that Koji had used in elog 8562, and put them in the new ASC filter banks (and turned them off in the SUS-PRM_ASCYAW bank).  I also moved all the old scripts that were in .../scripts/ASC to an OLD subdirectory (the most recent edit is from 2009 sometime).  I then copied over the up and down scripts that Koji had written for his ASC test into the ..../scripts/ASC directory, and modified them to work with my new channels. 

I then tried locking, and wasn't very successful.  Actually, my best lock, ~4 minutes, including tweaking up the PRM alignment, was when the ASC path was off (even though I thought it was on).  After discovering my mistake, I tried locking for another hour or so, but haven't really gotten anywhere.  The lock stretches I'm getting are rarely long enough for me to get to the terminal and run my up script, and the maybe ~6 or 7 times I've been able to run it, I haven't converged toward finding a good gain value for the PRC yaw loop.  At some point, I redid the MICH alignment since it had drifted away a bit, but that didn't really help.

I think that one of the next things I might try is carrier-locking the PRMI, to find okay loop gain settings for the ASC path.  Since the QPD output is already normalized (I'd have to custom-make some electronics to make it non-normalized), I think the gain should be the same for both carrier and sideband lock cases.

_______________________________

Once I finally get a good, stable, PRMI sideband lock, I think I need to take the following measurements:

* CTRL and ERR spectra for MICH and PRCL

* TFs for MICH and PRCL loops

* Sensing matrix, including AS55, REFL11, REFL33, REFL55, POX and POY.

---->> Are there any others?

ELOG V3.1.3-